# Establishing validated standards for Home and Work location Detection

### v1

@@ 2 Results
2.1 Unveiling temporal patterns of home-work visits
We analyze the regularity of home and work visits by measuring how often and when individuals are recorded at their annotated locations. Figure 1a-b shows the probability of being at home or work by hour of the day. We find that home visits peak between midnight and 6 AM, while work visits are most frequent during typical business hours (between 9 AM and 4 PM).
Patterns of visits to the home locations are similar between weekdays and weekends (see Figure 1c-e). For the work location, we observe a drastic reduction in the fraction of individual visiting their work location during weekends (see Figure 1d-f).
Results are consistent across both datasets, with overall higher probabilities of being at home (and lower probabilities of being at work) found in D2 due to the COVID-19 pandemic containment policies in place during most of the period covered by the data (see Sec.SI-3).
However, while general trends exist, we also find significant variations across days and users, highlighting pronounced behavioral heterogeneities at the individual level. Figure 1c-f reports disaggregated visit patterns for single “user-day” pairs. Most users’ daily patterns can be grouped into six behavioral profiles based on home and three on workday visits (see Methods for details on how behavioral profiles are obtained).
Figure 1: Behavioral visit profiles to home and work locations a. Fraction of visits per hour of the day to the annotated home locations in dataset D1, for weekdays (black) and weekends (grey) across individuals. The shaded orange area highlights the period with the highest probability of being at home. b. Fraction of visits per hour of the day to the annotated work locations in dataset 1, for weekdays and weekends. The shaded blue area highlights the period with the highest probability of being at work during weekdays. c. User-day profiles by hours at home (orange), other locations (grey), and periods without location data (light-grey) during weekdays. d. Hours at work (blue), other location (grey), or without data (light-grey), per hour of the day across days with data for weekdays. e-f. Show user-day profiles during weekends for hours at home (orange) and work (blue) respectively. The user-day profiles are sorted by cluster, hour of the first home visit, total time spent at home, and hours without data. Clusters are separated by black vertical lines, and the fraction of days per cluster is noted on the x-axis.
With respect to home visits, most user days fall into one of three typical profiles: i) days when users mostly stay at home, ii) when they are rarely/never at home, and iii) when they are away from home during typical business hours (see Figure 1c,e). Similarly, work-behavior profiles are described by three main daily patterns: i) days when users spend business hours at work, ii) when they are rarely/never at work, and iii) when they spend night hours at work (Figure 1d,f). We find that the distribution of individuals across these profiles is generally independent of their socio-demographic characteristics, such as age, gender, and urbanization (see Sec.SI-5).
As the behavioral profiles we identified are based on single user-day activity data, an important question is whether individuals switch between these profiles or, rather, a single profile consistently represents them. We find that users visit a median of $5$ out of $6$ home clusters and $2$ out of $3$ work clusters (for both D1 and D2). We measure the individual inter-profile variability through the normalized entropy, a measure ranging from 00 (when all days of a user fall into a single profile) to $1$ (when all profiles are equally important) [38] (see Methods). On average, users exhibit a normalized entropy of $0.66\pm 0.01$ for home-profile changes between weekdays and $0.48\pm 0.01$ for weekends ($0.62\pm 0.01$ and $0.42\pm 0.01$ for D2 respectively), and $0.51\pm 0.01$ for work-profile changes on weekdays only ($0.40\pm 0.02$ for D2). These values are comparable to the ones obtained when randomly allocating profiles among user days, confirming that users do not have a single dominant behavioral profile (see Table si.1).
All in all, our findings suggest that focusing on the same profiles across users is unlikely to favour one specific socio-demographic group, thus supporting the choice of focusing on weekday daytime hours (9 am to 4 pm) for detecting work locations and nighttime hours (12 am to 6 am) for detecting home locations, regardless of weekdays and weekends. These time windows, combined with minimum data coverage requirements, form the foundation of the heuristics implemented by the HoWDe algorithm.
2.2 HoWDe: a Home and Work Detection algorithm
In this section, we introduce HoWDe, a dynamic home and work detection algorithm that can capture residential moves and job changes over time through a sliding window approach. A key feature of HoWDe is the identification of home and work locations based on the proportion of time spent at a given location, rather than the absolute amount of time as approached in previous literature [29, 28, 14, 15, 33, 34]. This enables HoWDe to accommodate users with varying amounts of missing data over time (see Figure 1c-f), through tunable parameters that set data-quality filters. We share an open-source implementation of the algorithm in PySpark [37].
2.2.1 HoWDe workflow
HoWDe takes as input a user’s pre-processed sequence of stops, where each stop is fully defined by a user-location unique ID and a start and end timestamp.
First, we perform a temporal aggregation of the stop sequences into hourly-bin sequences. In each hourly bin, we retain the location where the individual has spent most of the time. This aggregation limits sequences to a maximum of 24 locations per day, boosting computational efficiency and making HoWDe faster than previous approaches based on sliding windows, with only limited information loss [14, 15].
Second, we exclude days with insufficient hourly bins with data (this step is controlled by the tunable parameter $dhn$, see Table 2). This is done separately for work detection (where we consider business hours) and home detection (where we consider night hours, see Figure 2b). Hence, it is possible that one day is retained only for the home detection or for the work detection.
Third, for each day $t$, we calculate the fraction of hourly-bins spent at each location during nighttime (24 pm to 6 am for home detection), and typical business hours (9 am to 4 pm on weekdays, for work detection). To compute such a fraction, we use as a denominator the number of hourly bins with records (Figure 2.b). Unlike previous approaches that rely on absolute time spent at each location, considering fractions can better accommodate users with varying amounts of missing data over time.
Fourth, for each day $t$, we consider a window of size $\Delta_{T}+1$, covering the range included in $t\pm\Delta_{T}/2$ days (see Figure 2.c). This window acts as a memory component for the algorithm to better detect behavior in the period before and after a specific day. In each window, we average the fraction of hours a location was visited (see Figure 2c).
Thus, for each window, we obtain a collection of potential home and work locations. Among these, we retain only those that meet minimum requirements with respect to the proportion of visits within and across days. These requirements are controlled by parameters $dn_{H,W}$, $hf_{H,W}$, $df_{H,W}$ (see Table 2). If no location meets the requirements, no home (or work) is assigned for that specific window.
Finally, we sort the home and work collections in two different ways. For home, we sort based on the typical fraction of hourly bins the location is visited within a day. For work, we sort based on the typical fraction of days the location is visited within the window. If no work location can be selected based on the fraction of days, the sorting is based on the fraction of hourly bins. The top location from these two lists is then assigned as home (or work). By prioritizing the fraction of days over the fraction of hours for work detection, we are emphasizing work as a location defined by regular visits throughout the workweek (see Sec.SI-6).
In summary, the algorithm allows the control of two aspects. It enables the user to specify the criteria for determining when there is enough data to attempt home and work detection. Also, it enables to control of when a location can be considered as a candidate home or work location, all while accounting for the fact that users can have different levels of missing data.
Figure 2: Schematics of the HoWDe algorithm workflow for home/work location detection. a. Network of visited locations. The diameter represents the number of visits. In orange, we report the stops visited during night hours, and in blue are the ones visited during work hours. b. Sequences of locations visited hourly. The pie charts display the allocation of visits across locations during the night hours (orange) and work hours (blue) on each day. Note that on Saturday, 6/01, the night hours data was discarded since there were not enough hours with available data (exemplifying the filter tuned by the parameter $dhn$). In this minimal example, no location appears during both night and work hours; however, this may happen in real data. c. For each day, $t$ we consider a sliding window of $3$ consecutive days centered on $t$. For work detection, the sliding window excludes weekends. d. Aggregation step for the window centered on Friday, 5th. For each location, we compute the average fraction of time it was visited during night hours (orange) and during work hours (blue). Additionally, we compute the fraction of days the location is visited at least once during work hours. Locations are then sorted in descending order based on these fractions, and the top ones are selected as the estimated Home (L1) and Work (L2) locations on Friday.
2.3 Validation against ground truth data
@@
To assess how well the procedure described above can capture individuals’ home and work locations, we measure i) the accuracy of the algorithm as the fraction of home/work locations that are correctly identified, and ii) the fraction of not-detected users, measured as the number of users where a home/work location was not detected against the number of users for which a home/work location was reported in the ground truth data (see Methods for more details). We run the validation using the ground-truth datasets D1 and D2, described in Methods. In dataset D2, annotators reviewed complete stop sequences; therefore, the temporal dimension (changes of home and work locations) can not be considered. For this reason, for D2, we consider a single window covering the entire data period (see Sec.SI-7 for parameter details).
@@ 2.3.1 Algorithm performance across parameter choices
We find that sliding window length $\Delta_{T}$ and minimum home-time fraction $hf_{H}$ have the greatest impact on the accuracy of HoWDe’s home detection, while other parameters have a smaller effect (see Sec.SI-8).
@@
Both the accuracy of the home location detection and the fraction of non-detected homes increase as the minimum home-time fraction threshold $hf_{H}$ increases (Figure 3a,b). In dataset D1, the fraction of non-detected homes increases with the window size, but accuracy shows no substantial improvement beyond $\Delta_{T}=28$ days ($t\pm 2$ weeks). This suggests that a window of $4$ weeks offers a good balance between high accuracy and user retention for home detection.
@@
Figure 3: Home and Work validation across configurations a. Accuracy of the home location detection for datasets D1 (first column) and D2 (second column). Each marker represents different configurations of the minimum fraction of hours an individual must be at home per night. For dataset D1, we explore how the results change for each sliding window increase (x-axis), while for dataset D2, we focus on a single sliding window (comprising the entire data period). b. Fraction of not detected of the home location detection for dataset D1 (first column) and dataset D2 (second column), sharing with the panel a) the legend. c. Accuracy of the work location detection for datasets D1 (first column) and D2 (second column). Each marker represents different configurations of the minimum fraction of hours an individual must be at work within the typical business hours, and the minimum fraction of days an individual should visit the potential work across weekdays. For dataset D1, we explore how the results change for each sliding window increase, while for dataset D2, we keep the same sliding window size (comprising the entire period). d. Fraction of not detected of the work location detection for dataset D1 (first column) and dataset D2 (second column). The legend is shared with panel c.
@@
Table 1 reports the results by focusing on two configuration setups: the minimum configuration, which maximizes accuracy while minimizing the number of non-detected users, and the maximum configuration, which prioritizes maximum accuracy at the cost of higher user non-detection rates. We also report the results obtained by two state-of-the-art home/work detection algorithms: Atlas [28, 2] and TimeGeo [29] (see Methods for more details on these algorithms).
@@
The maximum configuration is reached for $\Delta_{T}=28$ days window and $90\%$ minimum home-time fraction (see Sec.SI-7 for parameter details). With this choice of parameters, the algorithm achieves an accuracy of $97.30\%\pm 0.39\%$ for D1, and $88.47\%\pm 1.13\%$ for D2, but comes at the cost of excluding nearly two out of every five users (see Table 1). The minimum configuration is obtained for parameters $\Delta_{T}=28$ days window and $50\%$ home frequency (see Sec.SI-7 for parameter details). Here, the algorithm achieves $92.94\%\pm 0.49\%$ accuracy for D1, and $72.34\%\pm 1.13\%$ for D2, while maintaining a fraction of non-detected of $4.29\%\pm 0.38\%$ for D1 and $10.44\%\pm 1.02\%$ for D2.
@@
All in all, we find that HoWDe achieves higher accuracy than previous heuristic methods [29, 2], partly because it excludes users with low-quality data (see Table 1).
@@
Shifting to work location detection, we find that the fraction of correctly detected work locations increases with the window size until $\Delta_{T}=42$days ($t\pm 3$ weeks), after which accuracy stabilizes while non-detection rates continue to rise (see Figure 3c,d).
@@
In the work location detection, two key parameters are the minimum fraction of hourly bins ${hf_{W}}$ and days ${df_{W}}$ required for a location to be selected as work. Examining the impact of these two parameters (see Figure 3c-d, different colors), we find that stricter thresholds enhance accuracy in D1 but increase non-detection rates. For a larger value of ${hf_{W}}=60\%$, increasing ${df_{W}}$ enables significant accuracy gains. In contrast, for ${hf_{W}}=40\%$, changes in ${df_{W}}$ have a weaker effect. We find similar results for D2, when changing the minimum fraction of days ${hf_{W}}$ and fraction of hours ${hf_{W}}$.
@@
HoWDe achieves an average work detection accuracy of $83.84\%\pm 0.80\%$ and $88.14\%\pm 0.83\%$ for the minimum and maximum configurations in dataset D1 (see Sec.SI-7 for parameter details), outperforming baseline methods (see Table 1). Similarly, for D2, HoWDe achieves high accuracy levels ($81.86\%\pm 2.13\%-74.43\%\pm 2.13\%$), but, also in this case, systematically lower than D1 (see Table 1). This difference in performances may be partly explained by the fact that dataset D2 was collected during the COVID-19 pandemic, when people did not follow traditional mobility patterns.
Accuracy HoWDe min HoWDe max Atlas TimeGeo
Home D1 $\mathbf{92.94\pm 0.49}$ $\mathbf{97.30\pm 0.39}$ $81.96\pm 0.55$ $79.28\pm 0.78$
D2 $\mathbf{72.34\pm 1.13}$ $\mathbf{88.47\pm 1.13}$ $70.77\pm 1.21$ $70.38\pm 1.12$
Work D1 $\mathbf{83.84\pm 0.80}$ $\mathbf{88.14\pm 0.83}$ $60.32\pm 0.80$ $39.51\pm 0.92$
D2 $\mathbf{76.43\pm 2.13}$ $\mathbf{81.86\pm 2.18}$ $69.71\pm 2.17$ $57.04\pm 2.40$
@@
Work D1 $20.64\pm 0.77$ $39.71\pm 0.93$ $3.45\pm 0.31$ $\mathbf{1.86\pm 0.23}$
D2 $34.40\pm 1.93$ $58.73\pm 1.87$ $\mathbf{0.0\pm 0.0}$ $0.67\pm 0.41$
Table 1: Evaluation of home and work detection across methods and datasets.
2.4 Robustness and biases across configurations
HoWDe’s accuracy comes at the cost of user retention. The exclusion of users may introduce biases toward specific demographic groups. To assess these biases, we measure accuracy and user retention for different groups based on age (self-reported), gender (self-reported), urbanization levels, and country GDP. Results are obtained for dataset D1, using the minimum and maximum parameter configurations (see Sec.SI-7).
The largest accuracy gaps appear in suburban areas and middle-income countries: home detection drops by $3.39$ points in suburban areas (compared to urban areas) and work detection by $9.13$ in middle-income countries (compared to high-income countries). Despite accuracy deviations, home detection accuracy remains within $90\%$ $99\%$ and work detection within $77\%$ $89\%$ across all lower-accuracy subgroups (see Sec. SI-9), aligning with the observed range for D1 and D2.
Nevertheless, even minor detection errors can propagate, compromising the overall reliability of the results. To address this potential issue, we examine how tuning the detection algorithm parameters impacts results in two downstream applications: (i) employment rate estimates in Europe and (ii) commuting pattern characteristics.

### v2

@@ 6 Results
Here, we evaluate the performance of HoWDe against ground truth data, explore the effect of parameter choices, and assess its robustness across demographic groups. Then, we demonstrate how HoWDe can be applied to estimate employment rates and commuting patterns.
6.1 Algorithm performance: accuracy, retention, and robustness
@@
To assess how well HoWDe can capture individuals’ home and work locations, we measure i) the detected accuracy of the algorithm as the fraction of home/work locations that are correctly identified, and ii) the fraction of not-detected users, measured as the number of users where a home/work location was not detected against the number of users for which a home/work location was reported in the ground truth data (see Sec. 3.3 for more details). We run the validation using the ground truth datasets D1 and D2, described in Methods. In dataset D2, annotators reviewed complete stop sequences; therefore, the temporal dimension (changes of home and work locations) can not be considered. For this reason, for D2, we consider a single window covering the entire data period (see Supplementary material Sec. S-9 for parameter details).
@@ 6.1.1 Algorithm performance across parameter choices
We evaluate how the choice of configuration parameters affects HoWDe’s performance. Among all parameters, the sliding-window length $\Delta_{T}$ and the minimum home-time fraction $f_{hours}$ have the strongest influence on home-detection accuracy, while other parameters exert more moderate effects. Temporal coverage filters, in particular, primarily determine the amount of usable data and should therefore be tuned according to the quality and completeness of the dataset at hand. A detailed analysis of these additional parameter configurations, including their dependence on the datasets’ sparsity, is provided in Supplementary Material S-10.
@@
Both the detected accuracy of the home location estimation and the fraction of non-detected homes increase as the minimum home-time fraction threshold $f_{hours,H}$ increases (Figure 3a,b). In dataset D1, the fraction of non-detected homes increases with the window size, but the detected accuracy shows no substantial improvement beyond $\Delta_{T}=28$ days ($t\pm 2$ weeks). This suggests that a window of $4$ weeks offers a good balance between high accuracy and user retention for home detection.
@@
Figure 3: Home and Work validation across configurations a. Detected accuracy of the home location detection for datasets D1 (first column) and D2 (second column). Each marker represents different configurations of the minimum fraction of hours an individual must be at home per night ($f_{hours,H}$). For dataset D1, we explore how the results change for each sliding window increase (x-axis), while for dataset D2, we focus on a single sliding window (comprising the entire data period). b. Fraction of not detected home locations for dataset D1 (first column) and dataset D2 (second column), sharing with the panel a) the legend. c. Detected accuracy of the work location detection for datasets D1 (first column) and D2 (second column). Each marker represents different configurations of the minimum fraction of hours an individual must be at work within the typical business hours ($f_{hours,W}$), and the minimum fraction of days an individual should visit the potential work across weekdays ($f_{days,W}$). For dataset D1, we explore how results change with increasing sliding window sizes, while for dataset D2, we keep the same sliding window size (covering the entire period). d. Fraction of not detected work locations for dataset D1 (first column) and dataset D2 (second column). The legend is shared with panel c.
@@
Table 2 reports the results by focusing on two configuration setups: the minimum configuration, which maximises detected accuracy while minimising the number of non-detected users, and the maximum configuration, which prioritises maximum detected accuracy at the cost of higher user non-detection rates. We also report the results obtained by two state-of-the-art home/work detection algorithms: Atlas [32, 3] and TimeGeo [33] (see Secs. 2 and 3.4 for algorithm and performance comparison details).
@@
The maximum configuration is reached for $\Delta_{T_{H}}=28$ days window and $90\%$ minimum home-time fraction (see Supplementary material Sec. S-9 for parameter details). With this choice of parameters, the algorithm achieves a detected accuracy of $97.30\%\pm 0.39\%$ for D1, and $88.47\%\pm 1.13\%$ for D2, but comes at the cost of excluding nearly two out of every five users (see Table 2). The minimum configuration is obtained for parameters $\Delta_{T_{H}}=28$ days window and $50\%$ home frequency (see Supplementary material Sec. S-9 for parameter details). Here, the algorithm achieves $92.94\%\pm 0.49\%$ detected accuracy for D1, and $72.34\%\pm 1.13\%$ for D2, while maintaining a fraction of non-detected of $4.29\%\pm 0.38\%$ for D1 and $10.44\%\pm 1.02\%$ for D2.
@@
All in all, we find that HoWDe achieves higher detected accuracy than previous heuristic methods [33, 3], partly because it excludes users with low-quality data (see Table 2).
@@
Shifting to work location detection, we find that the fraction of correctly detected work locations increases with the window size until $\Delta_{T_{W}}=42$ days ($t\pm 3$ weeks), after which detected accuracy stabilizes while non-detection rates continue to rise (see Figure 3c,d).
@@
In the work location detection, two key parameters are the minimum fraction of hourly bins $f_{hours,W}$ and days $f_{days,W}$} required for a location to be selected as work. Examining the impact of these two parameters (see Figure 3c-d, different colors), we find that stricter thresholds enhance detected accuracy in D1 but increase non-detection rates. For a larger value of ${f_{hours,W}}=60\%$, increasing $f_{days,W}$} enables significant accuracy gains. In contrast, for ${f_{hours,W}}=40\%$, changes in $f_{days,W}$} have a weaker effect. We find similar results for D2, when changing the minimum fraction of days $f_{hours,W}$} and fraction of hours $f_{hours,W}$}.
@@
HoWDe achieves an average work detection accuracy of $83.84\%\pm 0.80\%$ and $88.14\%\pm 0.83\%$ for the minimum and maximum configurations in dataset D1 (see Supplementary material Sec. S-7 for parameter details), outperforming baseline methods (see Table 2). Similarly, for D2, HoWDe achieves high detected accuracy levels ($81.86\%\pm 2.13\%-74.43\%\pm 2.13\%$), but, also in this case, systematically lower than D1 (see Table 2). This difference in performance may be partly explained by the fact that dataset D2 was collected during the COVID-19 pandemic, when people did not follow traditional mobility patterns.
While Atlas does not explicitly estimate work locations, our adaptation (see Sec. 2) outperforms TimeGeo. This highlights that methods based solely on distance and visit counts, such as TimeGeo, perform substantially worse than those incorporating visit duration, including our Atlas adaptation and HoWDe. By accounting for missing data within individuals, HoWDe achieves the highest overall detected accuracy.
Detected accuracy HoWDe min HoWDe max Atlas TimeGeo
Home D1 $\mathbf{92.94\pm 0.49}$ $\mathbf{97.30\pm 0.39}$ $82.92\pm 0.66$ $81.44\pm 0.70$
D2 $\mathbf{72.34\pm 1.13}$ $\mathbf{88.47\pm 1.13}$ $70.62\pm 1.06$ $70.54\pm 1.15$
Work D1 $\mathbf{83.84\pm 0.80}$ $\mathbf{88.14\pm 0.83}$ $62.20\pm 0.92$ $37.51\pm 0.92$
D2 $\mathbf{76.43\pm 2.13}$ $\mathbf{81.86\pm 2.18}$ $71.74\pm 2.17$ $56.93\pm 2.35$
@@
Work D1 $20.64\pm 0.77$ $39.71\pm 0.93$ $\mathbf{0.0\pm 0.0}$ $0.80\pm 0.16$
D2 $34.40\pm 1.93$ $58.73\pm 1.87$ $\mathbf{0.0\pm 0.0}$ $0.68\pm 0.36$
Table 2: Performance comparison of home and work detection methods. Detected accuracy (top) and fraction not detected (bottom) across datasets D1 and D2. Bold highlights indicate the best method per metric—highest accuracy or lowest missing rate. While Atlas does not explicitly detect work locations, we adapt its home location algorithm for this purpose (see Section 2).
6.1.2 Algorithm robustness across demographic groups
HoWDe’s accuracy comes at the cost of user retention. The exclusion of users may introduce biases toward specific demographic groups. To assess these biases, we measure detected accuracy and user retention for different groups based on age (self-reported), gender (self-reported), urbanisation levels (estimated using the 2022 GHS settlement model grid at 1 km resolution [62]), and country GDP. Results are obtained for dataset D1, using the minimum and maximum parameter configurations (see Figure 4), and explored only for subgroups comprising at least 200 users.
The largest detected accuracy gaps appear in suburban areas and middle-income countries. In suburban settings, home detection accuracy drops by $3.39$ points compared to urban areas under the minimum configuration. However, user retention remains stable, suggesting that the decline in detected accuracy is not due to a disproportionate loss of suburban users.
In middle-income countries, home detection accuracy falls only under the maximum configuration (3.31-3.31 points), while work detection exhibits larger drops of $9.13$ and $8.06$ points under the maximum and minimum configurations, respectively. These declines occur without changes in user retention, suggesting that the reduced detected accuracy reflects differences in work visitation patterns rather than sample loss.
Across other demographic groups, we find significant differences in home detection. Home detected accuracy drops by $1.87$ points for males compared to females under the minimum configuration. For individuals aged 20–30, detected accuracy decreases by $2.40$ and $3.60$ points under the minimum and maximum configurations, respectively, relative to the 40–60 age group.
Despite accuracy deviations, home detection accuracy remains within $90\%-99\%$ and work detection within $77\%-89\%$ across all lower-accuracy subgroups (see Figure 4), aligning with the observed range for D1 and D2.
These findings emphasize that, while detection accuracy remains consistent across demographic groups and aligns with overall results, it is essential to validate sample representativeness when applying home and work detection algorithms.
Figure 4: Home and work detection accuracy and fraction of not detected locations across demographics for dataset D1. a. Detected accuracy of home location detection across demographic groups. Dark orange illustrates the results for the maximum configuration, while light orange illustrates the minimum. Error bars indicate bootstrapped error estimates (applies to all panels). b. Fraction of not detected home locations (configurations color coding is shared with panel a). c. Detected accuracy of work location detection across demographic groups. Dark blue illustrates the results for the maximum configuration, while light blue illustrates the minimum. d. Fraction of not detected work locations (configurations color coding is shared with panel c).
6.2 Impact of parameter sensitivity on real-world applications
Small detection errors can propagate across applications, potentially undermining the reliability of downstream results. To address this potential issue, we examine how tuning the detection algorithm parameters impacts results in two downstream applications: (i) employment rate estimates in Europe and (ii) commuting pattern characteristics.